How do I put this equation into standard form? Help me, please.
5x^2 - 4y^2 - 40x - 16y - 36 = 0

Answer by lwsshak3(11628)

How do I put this equation into standard form? Help me, please.
5x^2 - 4y^2 - 40x - 16y - 36 = 0
complete the square
5(x^2-8x+16)-4(y^2+4y+4)=36+80-16
5(x-4)^2-4(y+2)^2=100
(x-4)^2/20-(y+2)^2/25=1
This is an equation of a hyperbola with horizontal transverse axis:
center: (4,-2)
a^2=20
b^2=25
..
Standard form of equation for a hyperbola with horizontal transverse axis:
(x-h)^2/a^2-(y-k)^2/b^2=1, (h,k) being the (x,y) coordinates of the center